The same heat transfer into identical masses of different substances produces different temperature changes. Calculate the final temperature in degrees Celsius when 1.25 kcal of heat enters 1.25 kg of the following, originally at 15.0°C.

(a)water °C

(b)concrete°C

(c)steel °C

(d) mercury °C
Answer #1

Given, &= 1.25 kal= 5230丁 m=1.25kg sbeusfiz heat Capaun using formula, Q=mcPAT ve an Finå final JHembexatute of each Bubsremces , 5230 J = CI·25kg)1880 J/kc)CTf-15) ° Tf = 19.75C 8.54 Tf-15 29.88c= Tf-5 Tf = 44.88℃

Alright Dude, If that worked for you... dont forget to give THUMBS UP.(that will work for me!)

If I missed something feel free to leave a comment.

atleast before giving downvote.